Bill gave one-third of his jellybeans to a friend, ate two-thirds of the remaining jellybeans, and had eight jellybeans left over. How many jellybeans did Bill have originally?
Found 2 solutions by verde, ans_virgo4@hotmail.com:
Answer by verde(1)

if you work backwards it is easier to figure out the answer.
at the end he was left with 8 jellybeans.
before this he ate two thirds - this tells us that 8 equals one third, therefore he previosly had 24 jellybeans. (8*3=24)
if he gave one third to a friend then the remaining amount must be two thirds, therefore he began with 36 jellybeans. (24/2=12 12*3=36)

Answer by ans_virgo4@hotmail.com(2)

12*3=36 =3=36/12
=3=3
=3/3
=1ANS